OpenStudy (anonymous):

solve the system of equations, put answer in ordered pair form: 9x+2y=-46, -x+y=-1

OpenStudy (anonymous):

Multiply both sides of the second equation by \(-2\) to get \(2x-2y=2\). Add this new equation to the first one, you will have \(11x=-44 \implies x=-4\). Now plug x=-4 into the second equation, and you will find that \(4+y=-1 \implies y=-5\). The answer in ordered pair form is \((-4,-5)\).

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

\[4x=2y=-36\] is the I \[-x+y=-1\] is the II Operating II - I: \[-3x = -24\] \[x=8\] Going it in II: \[-(8)+y=-1\] \[y=7\] \[A={(8;7)}\]